OpenStudy (anonymous):

Hey guys... Here's a proof I am working on. I have to find out the untrue step in this proof. It's the 1+1 = 0 proof. Here are the steps: 1) 1 + 1 = 1 + √1 as we can take the value of √1 for 1 2) = 1 + √-1 * -1 because -1 * -1 is + 1 3) = 1 + √-1 * √ -1 now we can separate the multipliers and finally as we can denote √-1 with complex charecter i then 4) = 1 + i * i which can be written with squares like this 5) = 1 + i² as we know that i = √-1 and (√a)² is a with it we can write it like this now 6) = 1 + (√-1)² 7) = 1 + (-1) as + * - is -

OpenStudy (anonymous):

Here is where I think it is wrong: I don't think you can just separate the multipliers (now we can separate the multipliers and finally as we can denote √-1 with complex character) Is that right?

OpenStudy (anonymous):

if you need a better example, you can find the proof online by just typing in in 1+1 = 0 proof

OpenStudy (anonymous):

wer'z d rong step??

OpenStudy (anonymous):

That's what I need help figuring out!

OpenStudy (bahrom7893):

i don't think you can write (sqrt(-1))^2 = -1

OpenStudy (anonymous):

lol......i^2 =-1 ..so c iz right..

OpenStudy (anonymous):

i^2=-1(according to complex no.)

OpenStudy (anonymous):

i thinkk .ur stps r right ... wat makes u think dat u r goin wrong??

OpenStudy (anonymous):

because it's a trick question. There is suppose to be a wrong step in there.

OpenStudy (anonymous):

hold on, I figured it out. I'll upload the answer in a sec

OpenStudy (anonymous):

\[\int\limits(1/x)dx=\]

OpenStudy (anonymous):

1/x=u dv=dx

OpenStudy (anonymous):

hence , du=-1/x^2 v=x

OpenStudy (anonymous):

hence ... uv -\[\int\limits v.du\] =

OpenStudy (anonymous):

int(1/x)=x/x+int(1/x)

OpenStudy (anonymous):

hence 0=1

OpenStudy (anonymous):

therefore 1+1=0+0=0